# Discovery of the optical and radio counterpart to the fast X-ray transient EP 240315a

### v2

@@ Abstract
Fast X-ray Transients (FXTs) are extragalactic bursts of soft X-rays first identified years ago. Since then, nearly 40 events have been discovered, although almost all of these have been recovered from archival Chandra and XMM-Newton data. To date, optical sky surveys and follow-up searches have not revealed any multi-wavelength counterparts. The Einstein Probe, launched in January 2024, has started surveying the sky in the soft X-ray regime ( keV) and will rapidly increase the sample of FXTs discovered in real time. Here, we report the first discovery of both an optical and radio counterpart to an FXT, the fourth source publicly released by the Einstein Probe. We discovered a fast-fading optical transient within the 3 arcmin localisation radius of EP 240315a with the all-sky optical survey ATLAS, and our follow-up Gemini spectrum provides a redshift, . Furthermore, we uncovered a radio counterpart in the S-band (3.0 GHz) with the MeerKAT radio interferometer. The optical (rest-frame UV) and radio luminosities indicate the FXT most likely originates from either a long gamma-ray burst or a relativistic tidal disruption event. This may be a fortuitous early mission detection by the Einstein Probe or may signpost a mode of discovery for high-redshift, high-energy transients through soft X-ray surveys, combined with locating multi-wavelength counterparts.
@@ 1 Introduction
Alp & Larsson (2020) reported 12 FXTs from XMM-Newton and from inference of potential hosts they interpret the FXTs as emission from shock breakout in Wolf-Rayet stars within a dense circumstellar medium or (favoured in two cases) red supergiant progenitors. Eappachen et al. (2024) showed that seven of these have plausible host galaxies with spectroscopic redshifts , with one being a likely Galactic flare star. They proposed one FXT (XRT 110621) is consistent with being a supernova shock breakout (SBO) but the spectroscopic redshifts of the others showed that their peak X-ray luminosities were above that deemed feasible for supernova SBOs. Eappachen et al. (2024) also searched for contemporaneous optical counterparts in the Pan-STARRS and ATLAS wide-field surveys but found none. The detection limits range in depth (from to ; AB mags), and the delay between the bursts and the observations range from days. The most stringent limit on any contemporaneous optical emission remains the serendipitous observation of the location of CDF-S XT 1 with the Very Large Telescope (VLT) just 80 minutes after the burst (Bauer et al., 2017). With this observation, no associated optical counterpart – or host galaxy – was detected, down to a limiting -band magnitude, AB mag.
@@
In this Letter, we report the discovery of the optical and radio transient associated with EP 240315a, the first time multi-wavelength counterparts of an extragalactic FXT have been recorded. Throughout this paper we assume CDM cosmology with a Hubble constant, km s-1 Mpc-1, and (Planck Collaboration et al., 2016). We also assume a line-of-sight Milky Way extinction of AB mag, which corresponds to AB mag (Schlafly & Finkbeiner, 2011).
@@ 2.1 Discovery of the optical counterpart with ATLAS
The Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al., 2018a) is a quadruple 0.5-m telescope system, operating a wide-field all-sky survey. ATLAS continually surveys the sky, typically four times in 24 hr when all four units are operating normally, and we promptly process the data to search for extragalactic transients (Smith et al., 2020). During its normal survey operations, ATLAS observed the localisation region of EP 240315a at MJD 60384.894,111Here (and for all other optical observations), we quote the epoch of observation as the midpoint of the exposure. corresponding to hr (note that the first of the four 30 s exposures was obtained at MJD 60384.88673, or hr). Recall is the time of the detection from the Einstein Probe (MJD 60384.84079; Zhang et al., 2024c).
@@ 2.2 Optical photometric follow-up
The 2-m Liverpool Telescope (LT; Steele et al., 2004) was triggered under the program PL24A28 (PI: S. Srivastav). Images were obtained in -bands commencing on MJD 60385.847, corresponding to d. While the observing conditions were poor and the optical counterpart was not detected, our upper limit from non-detections confirmed the rapidly fading nature of AT 2024eju (see Srivastav et al., 2024a). Another set of -band images were obtained the next night, starting on MJD 60386.933 in better conditions. However, given the continuing rapid fade, AT 2024eju was only detected in -band. The derived magnitudes and upper limits from the LT images were estimated using the python-based Photometry Sans Frustration (PSF) code222https://github.com/mnicholl/photometry-sans-frustration (Nicholl et al., 2023).
@@
Finally, we obtained an epoch of late-time -band imaging with the Gemini-North/GMOS-N instrument, under the program ID GN-2024A-Q-221 (PI: M. Huber), commencing on MJD 60403.257 ( d). These observations were reduced using the DRAGONS pipeline (Labrie et al., 2023, 2023), and following standard recipes. AT 2024eju was not detected in these deep stacked images, and the upper limits were again derived using the PSF code.
@@ 2.4 Radio observations
Upon the discovery of a radio counterpart with the MeerKAT radio telescope, we obtained a rapid response time request with the enhanced-Multi-Element Radio Linked Interferometer Network (e-MERLIN, DD17004; PI: L. Rhodes). e-MERLIN is a UK-based radio interferometer with a maximum baseline of 217 km, and seven dishes spanning m in diameter. The facility can observe at L-, C- and K-band. Given the improved phase stability and sensitivity, we requested that our observation be made at C-band with a central frequency of 5.08 GHz and a bandwidth of 0.51 GHz. We obtained two observations; the first commenced on MJD 60389.736 ( d) and finished on MJD 60391.083 ( d), while the second started on MJD 60396.708 ( d) and finished on MJD 60398.042 ( d), each with a break in the middle whilst the target was below the horizon. The observations consisted of a series of six-minute scans of the target field, followed by two minutes on the phase calibrator (0933-0819). The target-phase cal loops were book-ended by visits to the flux and bandpass calibrators (J1331+3030 and J1407+2827, respectively). The data were flagged, calibrated and imaged using the e-MERLIN pipeline333https://github.com/e-merlin/eMERLIN_CASA_pipeline (Moldon, 2021). Using a uniform image weighting, we do not find any radio emission at either epoch, with 3 upper limits of Jy beam-1 and Jy beam-1, respectively. However, combining the observations reduced the rms noise to Jy beam-1, enabling us to extract a significant detection of Jy beam-1.
@@ 3.1 The early optical and radio fluxes
The optical discovery epoch (the cyan diamond data point in Figure 2) was obtained s (rest frame) after the EP-WXT stopped detecting the initial X-ray emission. It faded by magnitudes within rest-frame days ( to ; AB mags). The combination of proximity in time of the initial ATLAS detection and the X-ray counterpart, followed by such a rapid decay could indicate that the earliest optical emission is from the same emitting region and mechanism as the X-ray burst. Similar behaviour has been observed in some long GRBs where large field-of-view optical facilities have obtained simultaneous optical and gamma-ray detections (e.g., Vestrand et al., 2005; Racusin et al., 2008). It is important to distinguish between the prompt and afterglow emission in the optical data to create the most accurate picture of the early-time emission from this system for future modelling efforts.
@@ 3.2 What is the nature of EP 240315a/AT 2024eju?
GRBs are identifiable through their highly variable prompt gamma-ray emission, followed by a smoothly evolving synchrotron afterglow, produced as a highly relativistic jet collides with the circum-burst environment. GRB 240315C, detected by the Neil Gehrels Swift Observatory – Burst Alert Telescope (Swift/BAT) and Konus Wind (KW) instruments (DeLaunay et al., 2024; Svinkin et al., 2024a) was temporally coincident with EP 240315a. The BAT signal began at seconds and was detected at keV. The KW signal was detected from seconds at keV. The GRB detections lasted and 47 seconds in the BAT and KW data, respectively. The IPN triangulation of GRB 240315C places EP 240315a just within the annulus of 26.7∘ width (Svinkin et al., 2024b). Svinkin et al. (2024a) note that the KW detection had an elevated background due to increased solar flare activity. It is rare, but not completely unprecedented, to have a soft X-ray signal before the GRB itself (see e.g., Murakami et al., 1991; Piro et al., 2005), although in the case of EP 240315a and GRB 240315C the X-ray duration is significantly longer than the two previous cases. It is likely the two signals are related but further investigation of the high energy data is required.
@@
A short GRB (usually shorter than two seconds) interpretation of EP 240315a requires a BNS merger to have occurred Gyr after the Big Bang (derived from ), assuming standard cosmological parameters, as adopted in Section 1. Canonically dominant evolutionary channels involve an initially tight binary of massive OB stars which undergo two CCSNe (e.g., Tauris et al., 2017) and crucial common-envelope evolution before coalescence via gravitational inspiral. Binary evolution simulations find a range of peak delay time distributions from Gyr in rapid population synthesis (Belczynski et al., 2018; Chruslinska et al., 2018; Vigna-Gómez et al., 2018) to as low as 10 Myr in the BPASS detailed stellar evolution models (Eldridge et al., 2019). Additionally, the cosmic star-formation rate (and therefore the merger rate) is suppressed above (Madau & Dickinson, 2014; Mapelli & Giacobbo, 2018), but many models are consistent with significant BNS merger rates at (Santoliquido et al., 2021). Consequently, we cannot rule out a BNS origin for EP 240315a on the grounds of stellar evolution and inspiral time alone. However, we strongly disfavour the BNS merger/short GRB scenario through the comparison of radio luminosities and timescales. A sample of radio-detected short GRB afterglows is shown in the left-hand panel of Figure 4. Their luminosities are around two orders of magnitude lower than AT 2024eju, making a short GRB an unlikely origin of EP 240315a/AT 2024eju.
@@
Unlike short GRBs/BNS mergers, long GRBs (usually longer than two seconds), produced by collapsing massive stars,444We acknowledge the growing evidence for a population of merger-GRB events, including GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Gompertz et al., 2023) and GRB 230307A (Gillanders et al., 2023; Sun et al., 2023; Levan et al., 2024a; Yang et al., 2024) that lie within the long GRB s parameter space. However, here we are referring to the traditional progenitor system picture, where short GRBs are produced by compact object mergers and long GRB events are produced by massive star core-collapse. occupy a similar region of transient luminosity parameter space. In all wavebands, their afterglow component is more luminous than their short GRB analogues, likely due (at least in part) to higher kinetic energies in the jets of long GRBs compared to short GRBs (Fong et al., 2015; Aksulu et al., 2022). This is demonstrated best in the left-hand panel of Figure 4. The long GRB radio counterparts are approximately two orders of magnitude more luminous than short GRBs. As such, the long GRB radio population is far more consistent with the position of AT 2024eju in the radio luminosity parameter space. A similar conclusion can be reached regarding the position of AT 2024eju in optical luminosity parameter space. The right-hand panel of Figure 4 shows a sample of optically-detected, high-redshift (; therefore rest-frame UV) GRB afterglows, alongside our photometric measurements of AT 2024eju. Their luminosities and evolution are extremely consistent.
@@ 4 Summary and conclusions
The optical counterpart, AT 2024eju, was detected as a host-less transient by ATLAS during its routine all-sky survey operations just 0.054 days after the X-ray signal recorded by the Einstein Probe. We recorded a non-detection 1.7 days prior to discovery, and constrained its rapid fade, with it decaying by magnitudes in 19 hours post-discovery (see Figure 1).
The radio counterpart to EP 240315a was discovered by the MeerKAT radio telescope 1.86 days after the X-ray signal was recorded, and has been shown to originate from optically thick synchrotron emission by follow-up complementary e-MERLIN observations.
@@
The nature of this FXT indicates the Einstein Probe will uncover a range of high-energy transient phenomena in both the low- and high-redshift Universe.
@@
| All optical data will be included upon manuscript publication. | |||||
@@
| All radio data will be included upon manuscript publication. | |||||

### v3

@@ Abstract
Fast X-ray Transients (FXTs) are extragalactic bursts of soft X-rays first identified years ago. Since then, nearly 40 events have been discovered, although almost all of these have been recovered from archival Chandra and XMM-Newton data. To date, optical sky surveys and follow-up searches have not revealed any multi-wavelength counterparts. The Einstein Probe, launched in January 2024, has started surveying the sky in the soft X-ray regime ( keV) and will rapidly increase the sample of FXTs discovered in real time. Here, we report the first discovery of both an optical and radio counterpart to a distant FXT, the fourth source publicly released by the Einstein Probe. We discovered a fast-fading optical transient within the 3 arcmin localisation radius of EP 240315a with the all-sky optical survey ATLAS, and our follow-up Gemini spectrum provides a redshift, . Furthermore, we uncovered a radio counterpart in the S-band (3.0 GHz) with the MeerKAT radio interferometer. The optical (rest-frame UV) and radio luminosities indicate the FXT most likely originates from either a long gamma-ray burst or a relativistic tidal disruption event. This may be a fortuitous early mission detection by the Einstein Probe or may signpost a mode of discovery for high-redshift, high-energy transients through soft X-ray surveys, combined with locating multi-wavelength counterparts.
@@ 1 Introduction
Alp & Larsson (2020) reported 12 FXTs from XMM-Newton and from inference of potential hosts they interpret the FXTs as emission from shock breakout in Wolf-Rayet stars within a dense circumstellar medium or (favoured in two cases) red supergiant progenitors. Eappachen et al. (2024) showed that seven of these have plausible host galaxies with spectroscopic redshifts , with one being a likely Galactic flare star. They proposed one FXT (XRT 110621) is consistent with being a supernova shock breakout (SBO) but the spectroscopic redshifts of the others showed that their peak X-ray luminosities were above that deemed feasible for supernova SBOs. Soderberg et al. (2008) report an X-ray detection which they associate with the SBO from SN 2008D, at a distance of 27 Mpc (see also Chevalier & Fransson, 2008; Mazzali et al., 2008; Modjaz et al., 2009). Eappachen et al. (2024) also searched for contemporaneous optical counterparts in the Pan-STARRS and ATLAS wide-field surveys but found none. The detection limits range in depth (from to ; AB mags), and the delay between the bursts and the observations range from days. The most stringent limit on any contemporaneous optical emission remains the serendipitous observation of the location of CDF-S XT 1 with the Very Large Telescope (VLT) just 80 minutes after the burst (Bauer et al., 2017). With this observation, no associated optical counterpart – or host galaxy – was detected, down to a limiting -band magnitude, AB mag.
@@
In this Letter, we report the discovery of the optical and radio transient associated with EP 240315a, the first time multi-wavelength counterparts of a ‘distant’ ( Mpc) extragalactic FXT have been recorded. Throughout this paper we assume CDM cosmology with a Hubble constant, km s-1 Mpc-1, and (Planck Collaboration et al., 2016). We also assume a line-of-sight Milky Way extinction of AB mag, which corresponds to AB mag (Schlafly & Finkbeiner, 2011).
@@ 2.1 Discovery of the optical counterpart with ATLAS
The Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al., 2018a) is a quadruple 0.5-m telescope system, operating a wide-field all-sky survey. ATLAS continually surveys the sky, typically four times in 24 hr when all four units are operating normally, and we promptly process the data to search for extragalactic transients (Smith et al., 2020). During its normal survey operations, ATLAS observed the localisation region of EP 240315a at MJD 60384.894,111Here (and for all other optical imaging observations), we quote the epoch of observation as the midpoint of the exposure. corresponding to hr (note that the first of the four 30 s exposures was obtained at MJD 60384.88673, or hr). Recall is the time of the detection from the Einstein Probe (MJD 60384.84079; Zhang et al., 2024c).
@@ 2.2 Optical photometric follow-up
The 2-m Liverpool Telescope (LT; Steele et al., 2004) was triggered under the program PL24A28 (PI: S. Srivastav). Images were obtained in -bands commencing on MJD 60385.848, corresponding to d. While the observing conditions were poor and the optical counterpart was not detected, our upper limit from non-detections confirmed the rapidly fading nature of AT 2024eju (see Srivastav et al., 2024a). Another set of -band images were obtained the next night, at MJD 60386.946, in better conditions. However, given the continuing rapid fade, AT 2024eju was only detected in -band. The derived magnitudes and upper limits from the LT images were estimated using the python-based Photometry Sans Frustration (PSF) code222https://github.com/mnicholl/photometry-sans-frustration (Nicholl et al., 2023).
@@
Finally, we obtained an epoch of late-time -band imaging with the Gemini-North/GMOS-N instrument, under the program ID GN-2024A-Q-221 (PI: M. Huber), at MJD 60403.257 ( d). These observations were reduced using the DRAGONS pipeline (Labrie et al., 2023, 2023), and following standard recipes. AT 2024eju was not detected in these deep stacked images, and the upper limits were again derived using the PSF code.
@@ 2.4 Radio observations
Upon the discovery of a radio counterpart with the MeerKAT radio telescope, we obtained a rapid response time request with the enhanced-Multi-Element Radio Linked Interferometer Network (e-MERLIN, DD17004; PI: L. Rhodes). e-MERLIN is a UK-based radio interferometer with a maximum baseline of 217 km, and seven dishes spanning m in diameter. The facility can observe at L-, C- and K-band. Given the improved phase stability and sensitivity, we requested that our observation be made at C-band with a central frequency of 5.08 GHz and a bandwidth of 0.51 GHz. We obtained two observations; the first commenced on MJD 60389.736 ( d) and finished on MJD 60391.083 ( d), while the second started on MJD 60396.708 ( d) and finished on MJD 60398.042 ( d), each with a break in the middle whilst the target was below the horizon. The observations consisted of a series of six-minute scans of the target field, followed by two minutes on the phase calibrator (0933-0819). The target-phase cal loops were book-ended by visits to the flux and bandpass calibrators (J1331+3030 and J1407+2827, respectively). The data were flagged, calibrated and imaged using the e-MERLIN pipeline333https://github.com/e-merlin/eMERLIN_CASA_pipeline (Moldon, 2021). Using a uniform image weighting, we do not find any radio emission at either epoch, with 3 upper limits of Jy beam-1 and Jy beam-1, respectively. However, combining the observations reduced the rms noise to Jy beam-1,444The significant reduction in the rms noise of the concatenated observation is achieved because of our ability to recover otherwise-flagged data that are flagged out during the reduction of the individual images. enabling us to extract a significant detection of Jy beam-1.
@@ 3.1 The early optical and radio fluxes
The optical discovery epoch (the cyan diamond data point in Figure 2) was obtained s (rest frame) after the EP-WXT stopped detecting the initial X-ray emission. It then faded by magnitudes within rest-frame days ( to ; AB mags), corresponding to a temporal index of . The combination of proximity in time of the initial ATLAS detection and the X-ray counterpart, followed by such a rapid decay could indicate that the earliest optical emission is from the same emitting region and mechanism as the X-ray burst. Similar behaviour has been observed in some long GRBs where large field-of-view optical facilities have obtained simultaneous optical and gamma-ray detections (e.g., Vestrand et al., 2005; Racusin et al., 2008). It is important to distinguish between the prompt and afterglow emission in the optical data to create the most accurate picture of the early-time emission from this system for future modelling efforts. The afterglow component of the observed emission appears to flatten, or plateau, in the optical () bands.
@@ 3.2 What is the nature of EP 240315a/AT 2024eju?
GRBs are identifiable through their highly variable prompt gamma-ray emission, followed by a smoothly evolving synchrotron afterglow, produced as a highly relativistic jet collides with the circum-burst environment. GRB 240315C, detected by the Neil Gehrels Swift Observatory – Burst Alert Telescope (Swift/BAT) and Konus Wind (KW) instruments (DeLaunay et al., 2024; Svinkin et al., 2024a) was temporally coincident with EP 240315a. The BAT signal began at seconds and was detected at keV. The KW signal was detected from seconds at keV. The GRB detections lasted and 47 seconds in the BAT and KW data, respectively. The IPN triangulation of GRB 240315C places EP 240315a just within the annulus of 26.7∘ width (Svinkin et al., 2024b). Svinkin et al. (2024a) note that the KW detection had an elevated background due to increased solar flare activity. It is rare, but not completely unprecedented, to have a soft X-ray signal before the GRB itself (see e.g., Murakami et al., 1991; Piro et al., 2005), although in the case of EP 240315a and GRB 240315C the X-ray duration is significantly longer than the two previous cases. It is likely the two signals are related, but further investigation of the high energy data is required.
It is possible that EP 240315a and GRB 240315C fall within the class of ultra-long GRBs. Ultra-long GRBs are events whose prompt emission lasts as long as 10 000 seconds. Like ‘regular’ long GRBs, ultra-long GRBs have a large range of afterglow luminosities (see e.g., Levan et al., 2014, for multi-wavelength studies). As such, we cannot rule out the possibility that EP 240315a and GRB 240515C correspond to an ultra-long GRB event.
@@
A short GRB (events where the prompt flash of gamma-rays is usually shorter than two seconds)555This inferred duration is detector-dependent. interpretation of EP 240315a requires a BNS merger to have occurred Gyr after the Big Bang (derived from ), assuming standard cosmological parameters, as adopted in Section 1. Canonically dominant evolutionary channels involve an initially tight binary of massive OB stars which undergo two CCSNe (e.g., Tauris et al., 2017) and crucial common-envelope evolution before coalescence via gravitational inspiral. Binary evolution simulations find a range of peak delay time distributions from Gyr in rapid population synthesis (Belczynski et al., 2018; Chruslinska et al., 2018; Vigna-Gómez et al., 2018) to as low as 10 Myr in the BPASS detailed stellar evolution models (Eldridge et al., 2019). Additionally, the cosmic star-formation rate (and therefore the merger rate) is suppressed above (Madau & Dickinson, 2014; Mapelli & Giacobbo, 2018), but many models are consistent with significant BNS merger rates at (Santoliquido et al., 2021). Consequently, we cannot rule out a BNS origin for EP 240315a on the grounds of stellar evolution and inspiral time alone. However, we strongly disfavour the BNS merger/short GRB scenario through the comparison of radio luminosities and timescales. A sample of radio-detected short GRB afterglows is shown in the left-hand panel of Figure 4. Their luminosities are around two orders of magnitude lower than AT 2024eju, making a short GRB an unlikely origin of EP 240315a/AT 2024eju.
@@
Unlike short GRBs, long GRBs (events where the prompt flash of gamma-rays is usually longer than two seconds),5 produced by collapsing massive stars,666We acknowledge the growing evidence for a population of merger-GRB events, including GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Gompertz et al., 2023) and GRB 230307A (Gillanders et al., 2023; Sun et al., 2023; Levan et al., 2024a; Yang et al., 2024) that lie within the long GRB s parameter space, where is the time between the burst emitting 5 and 95 per cent of the detected counts. However, here we are referring to the traditional progenitor system picture, where short GRBs are produced by compact object mergers and long GRB events are produced by massive star core-collapse. occupy a similar region of transient luminosity parameter space. In all wavebands, their afterglow component is more luminous than their short GRB analogues, likely due (at least in part) to higher kinetic energies in the jets of long GRBs compared to short GRBs (Fong et al., 2015; Aksulu et al., 2022). This is demonstrated best in the left-hand panel of Figure 4. The long GRB radio counterparts are approximately two orders of magnitude more luminous than short GRBs. As such, the long GRB radio population is far more consistent with the position of AT 2024eju in the radio luminosity parameter space. A similar conclusion can be reached regarding the position of AT 2024eju in optical luminosity parameter space. The right-hand panel of Figure 4 shows a sample of optically-detected, high-redshift (; therefore rest-frame UV) GRB afterglows, alongside our photometric measurements of AT 2024eju. Their luminosities and evolution are extremely consistent.
@@
We rule out the possibility of AT 2024eju being an FBOT-like transient based on the mismatch between the early evolution of this FXT and that of AT 2018cow, the prototypical FBOT transient. While the peak bolometric luminosity of AT 2018cow roughly matches the early follow-up observations of AT 2024eju ( erg s-1; Prentice et al., 2018), the rise time is much slower; rest-frame days, versus rest-frame days for AT 2024eju. Additionally, the early radio lightcurve of AT 2018cow demonstrates a slow rise to maximum radio luminosity of rest-frame days (see e.g., Ho et al., 2019), which does not match the early, very luminous radio detection for EP 240315a. This FXT event evolves on a much more rapid timescale than AT 2018cow, both in the optical and radio, leading us to rule out FBOTs as an explanation for this transient.
@@ 4 Summary and conclusions
The optical counterpart, AT 2024eju, was detected as a host-less transient by ATLAS during its routine all-sky survey operations just 0.054 days (1.28 hours) after the X-ray signal recorded by the Einstein Probe. We recorded a non-detection 1.7 days prior to discovery, and constrained its rapid fade, with it decaying by magnitudes in 19 hours post-discovery (see Figure 1).
The radio counterpart to EP 240315a was discovered by the MeerKAT radio telescope 2.86 days after the X-ray signal was recorded, and has been shown to originate from optically thick synchrotron emission by follow-up complementary e-MERLIN observations.
@@
While it is difficult to quantify the future rates of such events, the discovery of EP 240315a so soon after the launch of the EP ( months) indicates that such events are probably not intrinsically rare. The soft X-ray regime that EP is optimised to explore is ideal for searching for high-redshift events that emit high-energy radiation, as redshifting will shift the peak of this emission from - to X-rays, making them more detectable to the EP, and other X-ray observatories. The nature of this FXT indicates the Einstein Probe will uncover a range of high-energy transient phenomena in both the low- and high-redshift Universe.
@@
| 60380.891 | ATLAS | 120 | |||
| 60383.145 | ATLAS | 120 | |||
| 60384.894 | ATLAS | 120 | |||
| 60385.673 | SLT | 1800 | |||
| 60385.848 | LT | 180 | |||
| 60385.851 | LT | 180 | |||
| 60385.853 | LT | 180 | |||
| 60386.144 | ATLAS | 120 | |||
| 60386.324 | PS | 300 | |||
| 60386.327 | PS | 300 | |||
| 60386.363 | PS | 700 | |||
| 60386.368 | PS | 700 | |||
| 60386.372 | PS | 700 | |||
| 60386.642 | LOT | 3000 | |||
| 60386.946 | LT | 2400 | |||
| 60386.976 | LT | 2400 | |||
| 60387.318 | PS | 2000 | |||
| 60387.346 | PS | 2000 | |||
| 60387.375 | PS | 1600 | |||
| 60388.307 | PS | 1600 | |||
| 60388.331 | PS | 2000 | |||
| 60388.356 | PS | 2000 | |||
| 60388.573 | SLT | 8700 | |||
| 60388.588 | LOT | 9000 | |||
| 60389.558 | LOT | 6000 | |||
| 60390.317 | PS | 2400 | |||
| 60390.347 | PS | 2400 | |||
| 60390.378 | PS | 2400 | |||
| 60403.257 | Gemini | 910 | |||
| 60403.276 | Gemini | 1170 | |||
@@
| 60387.703 | MeerKAT | 3.06 | |||
| 60390.407 | e-MERLIN | 5.01 | |||
| 60397.375 | e-MERLIN | 5.01 | |||